Johnny is Betty's father and is four times older than Betty. Five years ago, he was 7 times older than Betty. How old is Betty's father?
Answer by tiffany222(56)

Let the father's age be x
Let Betty's age be y
x = 4y
x - 5 = 7 (y-5)
4y - 5 = 7 (y-5)
4y - 5 = 7y - 35
-3y = -30
y = 10
x = 4y = 4(10) = 40
Final Answer:
Father's age = 40 years old